OpenStudy (anonymous):

f(x)=2sinxcosx, what is the primitive function?

OpenStudy (anonymous):

you have to find f'(x)

OpenStudy (anonymous):

That would be f'(x)=2cos^2(x)-2sin^2(x)

OpenStudy (anonymous):

I believe atleast, having such a hard time following productrules et cetera backwards.. :x

OpenStudy (anonymous):

\[\Large f'(x)=(2 \cdot \sin x \cdot \cos x)'\] take number 2 off... \[\Large f'(x)=2( \sin x \cdot \cos x)'\] and go on...

OpenStudy (anonymous):

http://www.sosmath.com/CBB/viewtopic.php?f=3&t=57694 the same problem I had before !

OpenStudy (anonymous):

I think you're supposed to rewrite the function into something more manageable.

OpenStudy (anonymous):

Perhaps by trigonometric identities or w/e, bah - clueless here!

OpenStudy (anonymous):

I don't understand you O_O

OpenStudy (anonymous):

is in the link what you're asking for or not ? ...(Google translate sucks !)

OpenStudy (anonymous):

Sorry, English is not my native language and I take math in Sweden - quite hard to "translate" the math into English, hehe.

OpenStudy (anonymous):

I cant put the thing you linked into context, hmm

OpenStudy (mimi_x3):

Primitive means antiderivative so integrate it. \[\int\limits2sinxcosx \] let u = cosx => du/-sinx

OpenStudy (anonymous):

@Mimi_x3 but how? ... then my task there was wrong , wan't it ? O_O

OpenStudy (anonymous):

I'm aware of that Mimi_x3 The thing is - I don't know how to do it! :D

OpenStudy (mimi_x3):

I remember when primitive means to integrate..

OpenStudy (mimi_x3):

not differentiate.

OpenStudy (anonymous):

Indeed it is the antiderivative.

OpenStudy (anonymous):

aufff... then I apologize , I learned it wrong :( ... thanks for explaining it @Mimi_x3

OpenStudy (mimi_x3):

Well, it's a u-substitution.

OpenStudy (mimi_x3):

Hm, but I'm not sure, lol.

OpenStudy (anonymous):

You're supposed to do it without substitution I believe, as the book I have explains substitution in the upcoming chapter.

OpenStudy (mimi_x3):

You can work backwards..when you find the derivative i think..

OpenStudy (anonymous):

I think what you basicly should do is use the product rule backwards, or rewrite it with the help of some trig identity.

OpenStudy (anonymous):

Indeed, it's very tiresome and confusing though, haha :(

OpenStudy (anonymous):

.. if it's integral ... then check this out ! ... http://www.wolframalpha.com/input/?i=2sinxcosx

OpenStudy (mimi_x3):

But she has not learnt u-substition.

OpenStudy (apoorvk):

PRIMITIVE = INTEGRAAAAAALLLL!!!!!

OpenStudy (anonymous):

I'm out !

OpenStudy (apoorvk):

write that as sin2x. then integrate using std. formula!!! should get ----> (-cos2x)/2 + constant

OpenStudy (anonymous):

I shall try, thanks everyone!